Shift Summary for Eve shift, 09.11.99

Patricia McBride (mcbride@fnpsph)
Sun, 12 Sep 1999 00:02:45 CDT

Shift Summary for Eve shift, 09.11.99

On Shift: P. McBride, Sasha Glasov

Also present: Ed, Sydney, JasonL, Julie, Ron, BobT, Peter Sh

Hours of beam delivered at all intensities: 8.25
Hours of beam delivered at nominal intensity: 8.25
Hours of E832 data (special runs not included): 4.75
Hours of special runs (laser scans, muon runs...): .30

Accounting of lost time: CsI sector 8 Low Voltage problems
========================================================================
Runs ended this shift:

14501 BEAM_E832 9.8 M events to tape
14502 MU_4E12 0.609 M muons at nominal settings for alignment
14503 MU_4E12 1.3 M muons for DC studies (3 NM4AN magnet settings)

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:

Julie, Ron and BobT made an access to repair the CsI Sector 8 low voltage
========================================================================
Overview of shift:

CsI LV power problems in Sector 8 were fixed during an access
Muon runs including special muon run for DC studies and magnet polarity flip
were done after the access followed by stable BEAM_E832 running.

========================================================================
Description of Shift:

run spill time comment
---------------------------------------------------------------------------
14501 15:30 Running smoothly

203 15:45 We notice a vertical strip missing from the top of
the laser illumination plot
We try resetting Etot, HCC but this does not help;
page Sydney who says he will come in to have a look

15:47 CSI low voltage alarm pops up; page Ed, page Julie
call for access; end run

16:15 Sydney arrives; Ron and Julie arrive; Ron and Julie go
on access to investigate the CsI LV; Sydney leaves

17:00 BobT arrives to join Ron and Julie on access

14502 18:00 Successful repair of CsI low voltage; start muon run

18:05 Call MCR to request beam; they say they will call
us back when they are ready

18:25 Operators visit to say there was a rad trip in NM3.
Most likely it was not real since there was no beam.
They believe it probably
happened when the beamstop was closed.
They will contact RSO to get the necessary approvals
to send beam to us.

29 18:35 Monitors reset; beam returns; running in muon mode

30-34 nominal muon run for alignment.
We tried to set the user latch to separate spills which
will have different analysis magnet settings for the
upcoming DC studies, however,
we could not view the tcl window for setting the
latches so we end the run and do a drt_fresh_start

14503 19:00 Muon Run for DC studies (3 settings)
current field ptkick
1-5 20 37.8 1.3
7-9 40 77.2 6.5
11-13 80 156.4 16.98

Set NM4AN magnet back to nominal setting (NEG polarity)

The dart_fresh_start did not help, so we did not
use the user latch. We rebooted the terminal and did
a dart_fresh_start again at the end of the run. The
Run Control windows are all working again.
(ST says that a simple logout would
probably have been sufficient.)

14504 19:30 BEAM_832 run started; stable running

138-139 DYC full (Laser crate), pipeline errors

23:00 We discover that the coffee grinder is broken;
Sasha uses the backup coffee grinder

0:00 Shift ends

========================================================================
Follow up Notes on Previous Shift

Beginning of shift summary for last evening read 9/2/99 instead of 9/10/99

========================================================================
Items to Follow up on Next Shift

If humidity alarms go off (humidity above 10% in the blockhouse)
call Leonard Nelson. His home number is 630-231-1416; pager
888-750-0702.